OpenStudy (anonymous):

which of the following functions has a maximum y-value of 2? a. y=2(x+1)^2+3 b. y=3(x+1)^2+2 c. y=-3(x+1)^2+2 d. y=-(x-2)^2+3

OpenStudy (anonymous):

\((x+1)^2\geq 0\) so \(-3(x+1)^2\leq 0\) so \(-3(x+1)^2+2\leq 2\)

OpenStudy (anonymous):

as usual the answer is C

OpenStudy (anonymous):

thanks for the help!